# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

### v1

@@ Abstract
Large Language Models (LLMs) have recently showcased strong capabilities in code-related tasks, yet their robustness in code comprehension and reasoning remains underexplored. In this paper, we present CodeCrash, a unified benchmark that evaluates LLM robustness under code structural and textual distraction perturbations, applied to two established benchmarks—CRUXEval and LiveCodeBench—across both input and output prediction tasks. We evaluate seventeen LLMs using direct and Chain-of-Thought inference to systematically analyze their robustness, identify primary reasons for performance degradation, and highlight failure modes. Our findings reveal the fragility of LLMs under structural noise and the inherent reliance on natural language cues, highlighting critical robustness issues of LLMs in code execution and understanding. Additionally, we examine three Large Reasoning Models (LRMs) and discover the severe vulnerability of self-reflective reasoning mechanisms that lead to reasoning collapse. CodeCrash provides a principled framework for stress-testing LLMs in code understanding, offering actionable directions for future evaluation and benchmarking. The code of CodeCrash and the robustness leaderboard are publicly available at https://donaldlamnl.github.io/CodeCrash/.
@@ 1 Introduction
Large Language Models (LLMs), exemplified by GPT-4 OpenAI et al. (2024), and the DeepSeek series DeepSeek-AI et al. (2025b, a), have recently exhibited remarkable performance across diverse code-related tasks, including code generation Zan et al. (2023); Xiao et al. (2024), program repair Fan et al. (2023); Liu et al. (2025), program testing Deng et al. (2023); Kang et al. (2023), and test case generation Steenhoek et al. (2025); Takerngsaksiri et al. (2025). Recent studies Wang et al. (2022); Mastropaolo et al. (2023); Chen et al. (2024a) have primarily evaluated LLM robustness under prompt-level perturbation, focusing on variations in task descriptions for code generation. While insightful for prompt engineering, these studies overlook robustness in deeper code comprehension tasks, such as understanding execution logic, structural reasoning, and program semantics.
In practical software development, LLMs are increasingly integrated into IDEs like GitHub Copilot [^1] https://github.com/features/copilot for tasks including code completion, automatic bug fixing, and test generation. Those integrations underscore the potential of LLMs to automate aspects of software development, improving code quality and developer productivity Zhang et al. (2023). Nevertheless, real-world production codebases are often complex and messy, frequently containing outdated comments, placeholder code, ambiguous identifiers, and non-functional segments, which introduce risk if LLMs overly depend on superficial cues rather than deep logical understanding.
CRUXEval (CRUXGu et al. (2024) introduces input and code execution tasks (i.e., output prediction) to evaluate LLMs’ comprehension and reasoning about code behavior. Building upon this foundation, we systematically evaluate LLM robustness under controlled code perturbations designed to challenge models’ reasoning capabilities beyond surface-level situations. Inspired by the program structure-consistent mutations proposed in CCTest Li et al. (2023), we extend this concept into a comprehensive robustness evaluation framework. Our unified benchmark, CodeCrash, systematically evaluates LLM robustness by applying two complementary perturbation families: I. Code Structural Perturbations, which introduce logically equivalent but structurally complicated code modifications, such as renaming entities, reformatting conditional expressions, and inserting non-functional “garbage” code segments; and II. Textual Distraction Perturbations, which insert misleading natural language (NL) cues that contradict actual program logic, including embedding misleading NL comments or statements and incorrect output hints.
@@
In this paper, we conduct extensive perturbation experiments on two established benchmarks: CRUX and LiveCodeBench (LCBJain et al. (2025). We evaluated the robustness of seventeen LLMs across distinct perturbations to assess their deep understanding of code behavior and functionality in input prediction and code execution tasks. Additionally, we apply Chain-of-Thought (CoT) prompting Kojima et al. (2022) to explore how explicit step-by-step reasoning enhances model performance and robustness in code comprehension. We further evaluate three state-of-the-art Large Reasoning Models (LRMs) with powerful internal reasoning processes, uncovering significant vulnerabilities arising from their self-reflection mechanisms. Our key findings reveal that:
  • LLMs struggle to strictly follow actual code behavior under structural perturbations, frequently distracted by irrelevant or non-functional segments.
  • LLMs inherently prioritize NL cues over actual program logic, heavily relying on identifiers, comments, and embedded messages, adopting superficial comprehension and shortcut reasoning.
  • Explicit step-by-step reasoning boosts performance and robustness; however, internal reasoning mechanisms may trigger severe overthinking, infinite self-reflection, and re-verification.
@@
  • We introduce a comprehensive perturbation framework CodeCrash encompassing structural complexity and textual distractions to evaluate the code-level robustness of LLMs.
  • We provide novel insights on evaluating and interpreting code reasoning in LLMs, revealing that they exhibit superficial NL dependence and demonstrating how plausible but incorrect hints trigger model collapsing.
@@ 2 Related Work
Researchers have significantly contributed to developing benchmarks for evaluating LLM capabilities on code-related tasks. Early benchmarks such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021a), APPS Hendrycks et al. (2021), and HumanEval+ Liu et al. (2023b) primarily focused on assessing code generation performance. More recent efforts have shifted toward practical software engineering and realistic coding scenarios. For instance, SWE-bench Jimenez et al. (2024), InfiBench Li et al. (2024), and LiveCodeBench Jain et al. (2025) evaluate LLMs on complex tasks like multi-file changes, Stack Overflow-style questions, and real-world programming problems.
Meanwhile, a growing number of works investigate LLMs in code execution. Early studies Austin et al. (2021b); Nye et al. (2021) indicated that LLMs face challenges in accurately executing code but can benefit from intermediate reasoning steps. Subsequent research has utilized various strategies to enhance execution performance, including execution-informed pretraining Liu et al. (2023a), CoT prompting with trace supervision Ni et al. (2024), and iterative instruction prompting Lyu et al. (2024). CruxEval Gu et al. (2024) formalized input and output prediction tasks specifically for evaluating LLMs’ code understanding and reasoning, while LiveCodeBench extended these assessments to realistic programs written by humans. REval Chen et al. (2024b) and CACP Hooda et al. (2024) uncovered inconsistencies and concept-level misunderstandings of LLMs in code through intermediate reasoning analysis in code execution. CodeScore Dong et al. (2024), LEVER Ni et al. (2023), and XCODEEval Khan et al. (2024) have leveraged code execution to improve model evaluation methodologies and enhance the performance of downstream tasks.
@@ 3.1 Benchmark Tasks and Datasets
We evaluate two benchmarks: CRUX Gu et al. (2024) and LCB Jain et al. (2025). CRUX contains 800 synthetic problems generated by Code-LLaMA-34B, each featuring concise Python snippets with generic function names f and a corresponding input-output pair for input prediction and code execution tasks. In contrast, LCB contains 479 real-world coding problems sourced from 85 human-submitted LeetCode solutions, focusing solely on code execution. While both datasets avoid excessive memory or runtime demands, LCB problems exhibit more complex structures, such as nested functions, decorators, and external libraries, requiring more computations.
@@ 3.2 Benchmark Execution Pipeline
Figure 1: Overview of CodeCrash pipeline. Code samples are first standardized into a clean vanilla version. Perturbations are visualized with yellow highlights. Each perturbed code is validated and executed before passing to LLMs.
Our benchmark follows a unified evaluation pipeline (Figure 1) across all perturbation types. Each original code snippet is initially processed via Abstract Syntax Tree (AST) parsing and unparsing to remove comments and standardize formatting, producing a clean baseline “vanilla” version (VAN). Subsequently, we apply a selected perturbation and regenerate the code using AST unparsing to maintain formatting consistency (except for comment-based perturbations). Each perturbed code snippet is executed with the provided inputs to ensure syntactic validity and functional correctness. The validated perturbed code is then provided to LLMs, and we record all execution results and evaluate them using the original evaluation methods from the dataset to compute the PASS@1 score.
@@ 3.3 Perturbation Strategies
We design two categories of perturbations to evaluate model robustness comprehensively: (1) Code Structural Perturbations, which modify code structure without affecting functionality, and (2) Textual Distraction Perturbations, which inject misleading NL cues and hints. Detailed examples and principles are provided in Appendix B.
@@ 3.3.1 Code Structural Perturbations
We insert syntactically valid but functionally irrelevant “garbage” code segments to challenge the models’ robustness against code structural distractions. Three specific types include: (1) repeated global variable declarations mirroring local identifiers, (2) non-executable conditional dead branches, and (3) uninvoked infinite recursive function definitions.
@@
We comprehensively aggregate all structural perturbations (REN, RTF, GBC) together and create a highly adversarial scenario, thoroughly challenging LLM robustness across multiple structural perturbation dimensions simultaneously.
@@ 3.3.2 Textual Distraction Perturbations
Similar to MDC perturbation, we inject misleading NL messages as executable print statements. Unlike comments that AST-based heuristics can easily filter out, print statements are logically irrelevant to functionality and should not be perceived as semantic guides. Allowing us to directly compare the behavior with MDC and distinguish models’ NL reliance in different embedding formats.
@@
We instruct GPT-4o to generate logically plausible but incorrect hints that contradict the actual program behavior and the ground truth. After verifying their incorrectness, they are strategically inserted into function definitions or return statements using AST parsing.
@@ 3.4 Experiment Settings
We evaluate 17 models of varying sizes and versions, including both open-source and closed-source LLMs: GPT family (GPT-4o-Mini, GPT-4o), Anthropic Claude (Claude-3.5-Haiku, 3.5-Sonnet), Google Gemini (Gemini-1.5-Flash, 2.0-Flash, 1.5-Pro), DeepSeek-V3, Alibaba Qwen (Qwen2.5-7B-Ins, 14B-Ins, 32B-Ins, 72B-Ins, 32B-Coder), and Meta LLaMA (LLaMA-3.1-8B-Ins, 70B-Ins, 405B-Ins, LLaMA-3.3-70B). Additionally, we assess three cutting-edge LRMs: o3-mini (low and high reasoning efforts), DeepSeek-R1, and QwQ-32B.
@@
Following the settings in LCB, we use nucleus sampling (temperature=0.2, top-p=0.95). We set the maximum tokens per response at 200 for direct inference and 2,000 for CoT inference, discarding excessively long outputs. For o3-mini, we set the maximum completion tokens at 20,000 to avoid verbose reasoning. Due to resource constraints, we generate three candidate answers for direct inference and one candidate for CoT inference per query. We adopt prompt formats similar CRUX, employing 2-shot prompting for direct inference and 1-shot CoT prompting (details provided in Appendix A). Primary evaluation uses PASS@1 accuracy Chen et al. (2021), reporting both absolute and relative degradation metrics.
4 Experimental Results
This section presents a comprehensive analysis of model robustness under the perturbation strategies introduced in §3.3. We assess the robustness of LLMs across different perturbation types, model scales, reasoning modes, and prediction tasks.
@@ 4.1 Impact on Output Prediction
We first evaluate LLM robustness in output prediction tasks. Table 1 summarizes the absolute PASS@1 degradation for all models under direct inference across both datasets, highlighting how each perturbation affects model performance.
Table 1: Absolute PASS@1 degradation of LLMs under all perturbation types for output prediction tasks across CRUX and LCB in direct inference (N=3) [TO RELATIVE SCORE].
@@
4.1.1 Robustness to Code Structural Perturbations
This section focuses on how structural perturbations impact model performance, examining reliance on superficial cues, logical reasoning capabilities, and sensitivity to irrelevant code.
Reliance on NL Cues.
Under the REN perturbation, we consistently observe performance degradation across GPT, Claude, DeepSeek, and Qwen models, with more significant drops on LCB ($8.20$pp 222pp stands for percentage point. average) than CRUX ($2.38$pp average). Top-performing models, such as Claude-3.5-Sonnet and DeepSeek-V3, suffer severe degradations exceeding $\mathbf{16}$pp on LCB. The substantial difference between the two datasets is attributed to the semantically meaningful function identifiers present in LCB, which are absent in CRUX. These results indicate that LLMs substantially rely on surface-level NL cues from identifiers to understand code and predict outputs.
Robust to Logical Reformatting.
LLMs show strong robustness against RTF perturbation, with minimal performance drop on both datasets (overall PASS@1 drop of only $1.63$pp). This robustness indicates LLMs can reliably reason over logically equivalent but syntactically complex expressions, demonstrating strong logical consistency against superficial structural modifications.
Severe Impact of Garbage Code.
Under the GBC perturbation, a notable performance degradation emerges in CRUX. The average degradation is consistently larger on CRUX ($7.99$pp) compared to LCB ($3.89$pp) across most models, which is a pattern opposite to other perturbations. This discrepancy can be attributed to the presence of meaningful identifiers in LCB, which helps models mitigate the distracting effects to infer the output. In contrast, the lack of meaningful identifiers in CRUX makes LLMs more susceptible to distraction by irrelevant and non-functional code segments. Under the fairness comparison environment in ALL perturbations, the performance drop on LCB ($15.50$pp) is again greater than that on CRUX ($13.17$pp), underscoring the significant impact of garbage code on LLMs’ reasoning and understanding.
Compounding Effects of Structural Perturbations.
Combining structural perturbations (ALL) results in the most severe degradation, averaging $\mathbf{14.04}$pp, which surpasses any individual perturbation effects and demonstrates an explicit compounding effect. Notably, high-performing models such as GPT-4o, Claude-3.5-Sonnet, Gemini-2.0-Flash, and DeepSeek-V3 suffer degradations exceeding $\mathbf{18}$pp on LCB, underscoring a serious vulnerability to multiple structural perturbations.
Key Findings: LLMs rely heavily on superficial semantic cues from identifiers, struggle to identify non-functional segments from essential code, and exhibit compounded vulnerability when multiple structural distractions are combined. This combination leads to significant degradation ($\mathbf{14.04}$pp on average) in output prediction performance.
4.1.2 Robustness to Textual Distraction Perturbations
We next investigate the influence of textual distractions on LLM robustness, focusing on the susceptibility to misleading information and incorrect hints.
Vulnerability to Misleading Information.
Misleading textual perturbations (MDC and MPS) cause substantial performance declines, averaging degradations of $13.47$pp and $13.28$pp, respectively—nearly as severe as the ALL perturbation. These results indicate that LLMs are highly susceptible to misleading NL messages, frequently prioritizing superficial textual information over the actual code logic. This behavior is further observed in MHC perturbation, causing an average degradation of $11.11$pp. The effect is notably higher on LCB ($17.83$pp) compared to CRUX ($7.09$pp), exceeding all other perturbations on LCB. This discrepancy indicates a difficulty-dependent vulnerability, wherein increased complexity exacerbates models’ reliance on plausible but misleading hints, diminishing their logical following and verifying abilities.
Consistency Patterns in Model Sensitivity.
We identify two prominent consistency patterns regarding sensitivity to misleading NL messages from Table 1. First, 16 of 17 models exhibit a clear dataset-level consistency: if the performance degradation caused by MDC on CRUX is greater than that on LCB, then the same relationship holds for MPS and vice versa. This pattern indicates that model sensitivity to misleading NL messages primarily depends on dataset characteristics rather than perturbation format, revealing an inherent NL-dependence property of LLMs in reasoning processing. Second, 14 of 17 models show a clear perturbation-method consistency: if the performance degradation caused by MDC on CRUX exceeds that on LCB, then the same holds for MPS and vice versa. This pattern suggests an internal cognitive preference for specific formats (comments vs. print statements). While the degradation magnitude varies with the code, the format-specific sensitivity is consistently preserved across datasets, reflecting the intrinsic properties of LLMs related to their reasoning architecture and training corpus.
Key Findings: Textual perturbations significantly degrade model performance, revealing an inherent dependence on embedded NL cues and a internal cognitive preference for specific message formats. Moreover, LLMs exhibit an explicit difficulty-dependent vulnerability, increasingly relying on plausible but incorrect hints over actual program logic in complex scenarios.

### v2

@@ Abstract
Large Language Models (LLMs) have recently demonstrated strong capabilities in code-related tasks, yet their robustness in code comprehension and reasoning remains insufficiently explored. We present CodeCrash, a comprehensive stress-testing benchmark comprising 1,279 questions from two established datasets, CruxEval and LiveCodeBench, designed to evaluate model reasoning reliability under non-standard coding environments. We systematically evaluate 17 LLMs across input and output prediction tasks using direct and Chain-of-Thought prompting approaches, revealing that LLMs are particularly vulnerable to disorganized code and overly reliant on natural language cues: aggregated structural perturbations result in over 14 percentage points (pp) of degradation, while textual perturbations cause a performance drop of over 11 pp. Moreover, self-reflective mechanisms in state-of-the-art reasoning models significantly increase token usage by 23×2-3$\times$2 - 3 ×, reduce output confidence, and even lead to catastrophic reasoning failures when faced with targeted perturbations—for instance, QwQ-32B generates over 12,000 redundant tokens under reasoning-level perturbations. CodeCrash provides a rigorous benchmark for evaluating robustness in code understanding, guiding future research toward more reliable and resilient LLMs in code reasoning. The benchmark code, perturbed datasets, and full leaderboard are publicly available at https://cuhk-arise.github.io/CodeCrash/.
@@ 1 Introduction
Large Language Models (LLMs), exemplified by GPT families OpenAI et al. (2024); Hurst et al. (2024) and the DeepSeek series DeepSeek-AI et al. (2025b, a), have recently exhibited remarkable performance across diverse code-related tasks, including code generation Zan et al. (2023); Liu et al. (2023b); Xiao et al. (2024), code completion Ding et al. (2023), program repair Fan et al. (2023); Liu et al. (2025), program testing Deng et al. (2023); Kang et al. (2023), and test case generation Steenhoek et al. (2025); Takerngsaksiri et al. (2025). These advancements have facilitated the incorporation of LLMs into Integrated Development Environments, such as GitHub Copilot GitHub (2022), OpenAI Codex Chen et al. (2021), and Tabnine Tabnine (2023), highlighting their potential to automate various aspects of software development Zhang et al. (2023); Pandey et al. (2024).
While the effectiveness of these applications heavily depends on the model’s ability to comprehend the intended program logic, underlying functionality, and developer intent Pan et al. (2024), real-world production codebases are often disorganized, with ambiguous identifiers Gao et al. (2023) and inconsistent comments Rong et al. (2025). Such non-standard artifacts can mislead models, leading to incorrect inferences and misinterpretations, compromising the model’s reliability in downstream tasks. Recent studies Wang et al. (2022); Mastropaolo et al. (2023); Chen et al. (2024a) have primarily evaluated LLM robustness under prompt-level perturbation, focusing on variations in task descriptions. While insightful for prompt engineering, these studies overlook robustness in deeper code comprehension tasks, such as reasoning about execution logic and program semantics.
To determine whether LLMs truly understand program logic, we adopt input and output prediction tasks introduced in CruxEval (CruxGu et al. (2024) as diagnostic methods for assessing model in code understanding and execution reasoning. Inspired by prior work on stress-testing Natural Language (NL) inference systems Naik et al. (2018), we propose a novel stress-testing strategy to evaluate the robustness of LLMs’ code comprehension capabilities under systematic code perturbations. First, building upon program structure-consistent mutations Li et al. (2023) and dead-loop poisoning Wan et al. (2022); Sun et al. (2023), we design Code Structural Perturbations, which involve logic-preserving transformations, such as renaming entities, reformatting conditional expressions, and inserting unreachable dead code. Second, motivated by inconsistent comments in real-world codebases Rong et al. (2025), we introduce Textual Distraction Perturbations, which inject misleading NL information into code to challenge the model’s ability to interpret program logic under conflicting cues.
@@
To extend our evaluation from synthetic programs to real-world codebases, we incorporate LiveCodeBench (LCBJain et al. (2025) alongside Crux as part of our benchmark suite. We evaluate seventeen LLMs across distinct perturbations in input and output prediction tasks. Additionally, we apply Chain-of-Thought (CoT) prompting Kojima et al. (2022) to explore how explicit step-by-step reasoning enhances model performance and robustness in code comprehension. Furthermore, we assess three state-of-the-art Large Reasoning Models (LRMs) with powerful internal reasoning processes, uncovering significant vulnerabilities arising from their self-reflection mechanisms. Our key findings reveal that:
  • LLMs perform unreliably under perturbations and heavily rely on embedded NL cues for reasoning and comprehension. Each textual distraction perturbation causes an average performance drop of over $11$pp 111pp stands for percentage point., while the aggregated structural perturbation also leads to up to $14$pp degradation.
  • CoT prompting improves model robustness, but cannot fully eliminate sensitivity to NL cues, resulting in a relative degradation of up to $14\%$ under contextual-level textual perturbation.
  • LRMs show superior robustness under most perturbations, but the self-reflection mechanism consumes 23×2-3$\times$2 - 3 × more reasoning tokens to resolve contradictions, even leading to reflection loops.
@@
  • We introduce a comprehensive perturbation framework, CodeCrash, to evaluate the code-level robustness of LLMs, which is easily applicable to other benchmarks and programming languages.
  • We conduct extensive experiments on 17 LLMs across two established datasets, Crux and LCB.
  • We provide novel insights into evaluating and interpreting code reasoning in LLMs, revealing the reliance on NL cues and demonstrating how reasoning-level hints can lead to model collapse.
@@ 2 Related Work
Researchers have significantly contributed to developing benchmarks for evaluating LLM capabilities on code-related tasks. Early benchmarks such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021), APPS Hendrycks et al. (2021), and HumanEval+ Liu et al. (2023b) primarily focused on assessing code generation performance. More recent efforts have shifted toward practical software engineering and realistic coding scenarios. For instance, SWE-bench Jimenez et al. (2024), InfiBench Li et al. (2024), and LCB Jain et al. (2025) evaluate LLMs on complex tasks like multi-file changes, Stack Overflow-style questions, and real-world programming problems.
Meanwhile, a growing number of works investigate LLMs in code execution (i.e. output prediction). Early studies Austin et al. (2021); Nye et al. (2021) indicated that LLMs face challenges in accurately executing code but can benefit from intermediate reasoning steps. Subsequent research has utilized various strategies to enhance execution performance, including execution-informed pretraining Liu et al. (2023a), CoT prompting with trace supervision Ni et al. (2024), and iterative instruction prompting Lyu et al. (2024). Crux Gu et al. (2024) formalized input and output prediction tasks specifically for evaluating LLMs’ code understanding and reasoning, while LCB extended these assessments to realistic programs written by humans. REVAL Chen et al. (2024b) and CACP Hooda et al. (2024) uncovered inconsistencies and concept-level misunderstandings of LLMs in code through intermediate reasoning analysis in code execution. CodeScore Dong et al. (2024), Lever Ni et al. (2023), and xCodeEval Khan et al. (2024) have leveraged code execution to improve model evaluation methodologies and enhance the performance of downstream tasks.
@@ 3.1 Benchmark Tasks and Datasets
We evaluate two benchmarks: Crux Gu et al. (2024) and LCB Jain et al. (2025). Crux contains 800 synthetic problems generated by Code-LLaMA-34B, each featuring concise Python snippets with generic function names f and a corresponding input-output pair for input and output prediction tasks. In contrast, LCB contains 479 real-world coding problems sourced from 85 human-submitted LeetCode solutions, focusing solely on output prediction. While both datasets avoid excessive memory or runtime demands, LCB problems exhibit more complex structures, such as nested functions, decorators, and external libraries, requiring more computation.
@@ 3.2 Benchmark Execution Pipeline
Our benchmark follows a unified evaluation pipeline (Figure 3) across all perturbations. Each original code snippet is initially processed via Abstract Syntax Tree (AST) parsing and unparsing to remove comments and standardize formatting, producing a clean baseline “vanilla” version (VAN). Subsequently, we apply a selected perturbation and regenerate the code using AST unparsing to maintain formatting consistency (except for comment-based perturbations). Each perturbed code snippet is executed with the provided inputs to ensure syntactic validity and functional correctness. The validated perturbed code is then provided to LLMs, and we record all execution results and evaluate them using the original evaluation methods from the dataset to compute the Pass@1 score.
@@ 3.3 Perturbation Strategies
We design two categories of perturbations to evaluate model robustness comprehensively: (1) Code Structural Perturbations, which modify code structure without affecting functionality, and (2) Textual Distraction Perturbations, which inject misleading NL cues and hints. Detailed examples and principles are provided in Appendix A.
@@ 3.3.1 Code Structural Perturbations
We insert syntactically valid but functionally irrelevant “garbage” code segments to challenge model robustness against code structural distractions. Three specific types include: (1) repeated global variable declarations mirroring local identifiers, (2) non-executable conditional dead branches, and (3) uninvoked infinite recursive function definitions.
@@
We comprehensively aggregate all structural perturbations (REN, RTF, GBC) together and create a highly adversarial scenario, thoroughly challenging LLM robustness across multiple structural perturbations simultaneously.
@@ 3.3.2 Textual Distraction Perturbations
Similar to MDC perturbation, we inject misleading NL messages as executable print statements. Unlike comments that AST-based heuristics can easily filter out, print statements are logically irrelevant to functionality and should not be perceived as semantic guides. Allowing us to directly compare the behavior with MDC and distinguish models’ NL reliance in different embedding formats. Since MDC and MPS perturbations inject shallow and contextually misleading information, we categorize them as contextual-level textual perturbations.
@@
We instruct GPT-4o to generate logically plausible but incorrect hints that contradict the actual program behavior and the ground truth. After verifying their incorrectness, they are strategically inserted into function definitions or return statements using AST parsing. Since these plausible hints challenge the model’s high-level inference ability, we classify MHC as a reasoning-level textual perturbation.
4 Experiments
This section presents a comprehensive analysis of model robustness under the perturbation strategies introduced in §3.3. We assess the reliability of LLMs across different perturbation types, model scales, reasoning modes, and prediction tasks.
@@ 4.1 Experiment Settings
We evaluate 17 models of varying sizes and versions, including both open-source and closed-source LLMs: GPT family (GPT-4o-Mini, GPT-4o) Hurst et al. (2024), Anthropic Claude (Claude-3.5-Haiku, 3.5-Sonnet) Anthropic (2024b, a), Google Gemini (Gemini-1.5-Flash, 2.0-Flash, 1.5-Pro) Pichai and Hassabis (2024); Pichai et al. (2024), DeepSeek-V3 DeepSeek-AI et al. (2025b), Alibaba Qwen (Qwen2.5-7B-Ins, 14B-Ins, 32B-Ins, 72B-Ins, 32B-Coder-Ins) Yang et al. (2025); Hui et al. (2024), and Meta LLaMA (LLaMA-3.1-8B-Ins, 70B-Ins, 405B-Ins, LLaMA-3.3-70B) Grattafiori et al. (2024). Additionally, we assess three cutting-edge LRMs: o3-mini OpenAI (2025), DeepSeek-R1 DeepSeek-AI et al. (2025a), and QwQ-32B Qwen Team (2025).
@@
Following LCB, we use nucleus sampling (temperature=0.2, top-p=0.95), with a maximum of 200 tokens for direct inference and 2,000 for CoT prompting, discarding excessively long outputs. For o3-mini, we raise the limit to 20,000 tokens to avoid verbose reasoning. Due to resource constraints, we generate three candidates for direct inference and one for CoT inference per query; we provide a result stability analysis under $N=3$ in Appendix B. Following Crux, we employ 2-shot prompting for direct inference and 1-shot for CoT inference (details provided in Appendix C). Furthermore, we adopt Pass@1 accuracy Chen et al. (2021) as the primary evaluation metric and scale all scores in $[0,1]$ to percentages by multiplying by $100$ for readability. All perturbed results are reported as either absolute ($\Delta_{\text{\;{pp}}}=\text{Perturbed}-\text{VAN}$) or relative ($\Delta_{\%}=\frac{\text{Perturbed}-\text{VAN}}{\text{VAN}}\times 100\%$) differences from the corresponding vanilla baseline. Additionally, we use percent signs (%) to indicate relative differences for clarity in all tables.
@@ 4.2 Impact on Output Prediction
We first evaluate LLM robustness in output prediction tasks. Table 1 summarizes the absolute Pass@1 degradation for all models under direct inference across both datasets. Our findings in this subsection include:
Key Findings: LLMs demonstrate an inherent and multi-layered reliance on natural language information, including identifiers, contextual messages, and reasoning hints, which ultimately undermines their reliability in code reasoning and comprehension. LLMs exhibit limited robustness to structurally disorganized code, struggling to differentiate irrelevant code from essential logic. Aggregating multiple structural perturbations exacerbates this vulnerability, significantly impairing their reasoning reliability. Larger scales, newer versions, and domain-specific fine-tuning LLMs substantially enhance robustness, highlighting improved code understanding and reasoning capabilities.
Table 1: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks on Crux and LCB using direct inference ($N=3$). VAN shows the Pass@1 accuracy (in %), others indicate absolute drops from VAN (in  pp). Darker red highlights represent more severe degradation. See [To Relative Version] for the relative version.
@@
4.2.1 Over-reliance on Natural Language Semantics
We categorize NL dependence into three hierarchical levels: surface-level, contextual-level, and reasoning-level dependence.
Surface-level NL Dependence (REN).
Under the REN perturbation, models suffer more significant performance degradation on LCB ($8.20$pp on average) compared to Crux ($2.38$pp) due to the presence of meaningful function names in the vanilla settings of LCB, which are absent in Crux. High-performing models like Claude-3.5-Sonnet and DeepSeek-V3 suffer severe degradations exceeding $\mathbf{16}$pp on LCB, highlighting a significant reliance on semantically meaningful identifiers.
Contextual-level NL Dependence (MDC and MPS).
Under contextual-level textual perturbations (MDC and MPS), LLMs exhibit severe performance drops ($13.47$pp and $13.28$pp, respectively). Notably, we identify two prominent consistency patterns: (1) dataset-level consistency [^2] If MDC degrades performance more on Crux than LCB, the same holds for MPS, and vice versa. ($16/17$ models), implying the model sensitivity to misleading comments and prints primarily depends on dataset characteristics and (2) perturbation-method consistency [^3] If MDC causes greater degradation on Crux than LCB, so does MPS, and vice versa. ($14/17$ models), indicating an inherent cognitive preference for NL messages from specific formats. These phenomena reveal that LLMs inherently depend on embedded NL context and exhibit format-specific cognitive biases.
Reasoning-level NL Dependence (MHC).
MHC perturbation also induces substantial degradation ($11.11$pp on average), notably exacerbated in complex datasets (LCB: $17.83$pp vs. Crux: $7.09$pp), which is a complexity-amplified effect. Advanced models like Claude-3.5-Haiku and Sonnet experience over $30$pp performance degradation, far exceeding their degradation under surface or contextual perturbations. These results illustrate that (i) LLMs readily accept plausible hint comments as authoritative reasoning shortcuts, and (ii) this over-trust intensifies when task complexity surpasses the model’s reasoning capacity, which we revisit under CoT prompting in §4.3.1.
4.2.2 Low Resistance to Disorganized Coding Environments
From RTF perturbation, LLMs can effectively reason over logically equivalent but syntactically complex expressions with minimal degradation ($1.63$pp). However, the highest performance drop emerges under the GBC compared to REN and RTF, with an average degradation of $6.46$pp across both datasets. Notably, GBC is the only perturbation where degradation is more severe on LCB ($3.89$pp vs. $7.99$pp on Crux) due to uneliminated meaningful identifiers. This trend reverses under the ALL perturbation, where all structural perturbations are combined, leading to a greater performance drop on LCB ($15.50$pp) than on Crux ($13.17$pp). Additionally, we observe a compounding effect of structural perturbations under ALL perturbation, resulting in the most significant overall degradation ($14.04$pp), which surpasses any individual code structural perturbation effects.